s ∗ ∗ ∗ ∗ ∗

March 27, 2009

The strange quark is the isospin=−1/2 second generation quark. The physical mass is at the centre of the range quoted in [?] and is only used as a coupling while the constituent mass is used kinematically. The s has mass 95 MeV and is stable. The s has spin 1/2, charge −1/3 and is in a colour triplet. The s is a quark. The particle is a constituent of the hadrons and the constituent mass is 500 MeV. The properties of the particle and its antiparticle are taken to be charge-conjugate to each other. The PDG code is 3. There are no decay modes for the s.

There is no mass generator for the s. There is no width generator for the s.

The particle was checked by Peter Richardson at 21:46:38 on the 20th of October 2007. The most recently changed decay mode was modified at 0:0:0 on the 0th of 0. The particle data was last modified at 16:27:46 on the 17th of September 2007.